# Portable Biomechanics Laboratory: Clinically Accessible Movement Analysis from a Handheld Smartphone

### v1

@@ Acknowledgements.
The outputs of $f_{\phi}$ corresponding to rotations (i.e., not the pelvis location) are passed through a tanh nonlinearity to limit it to $(-1,1)$ followed by scaling this range to match the biomechanical model joint limits. Joint angles, $\theta(t)$, are then passed through the forward kinematic equation above (Eq. 1) to obtain the 3D joint locations, $\mathbf{x}(t)$, at that time.
@@
We define the global reference frame $\{n\}$ for our model output. Since the PBL system tracks changes in phone orientation from the smartphone IMUs, we are able to relate the orientation of the camera frame $\{c\}$ to the orientation of the world frame at any point in time using $R_{nc}\in SO(3)$. We follow the convention presented in (63) for rotations where $R_{ab}$ represents frame $\{b\}$ relative to frame $\{a\}$.
@@
Each trial is represented as a unique implicit function. We jointly optimize the parameters of the implicit representations for N trials $\{\phi_{0},\phi_{1},...,\phi_{N-1}\}$, the body scaling, and site offset parameters $\beta$, to follow 2D and 3D keypoints extracted from video as well as orientation data recorded from the smartphone. This approach of jointly learning kinematics and body scaling and marker offset for the entire session has been called bilevel optimization (64). For example, a session consisting of 10 trials would jointly optimize 10 implicit functions $f_{\theta_{i}}$ and a single set of body scaling parameters $\beta$.
For each training step, we evaluated every implicit function with a batch of 300 time samples for each trial to extract the predicted joint angles at those time points $\hat{\theta}(t)$, which is performed in parallel across all trials. Using these poses and scaling parameters, we performed GPU accelerated forward kinematic passes of our model (Eq. 1) to predict 3D joint locations for that training batch, $\hat{\mathbf{x}}=\mathcal{M}(\hat{\theta}(t),\hat{\beta})$. Next, we describe the losses used during this optimization. To reduce notational clutter, we drop the explicit time and trial dependence in the loss definitions.
@@
We obtain a pure video estimate of the 3D keypoint locations in the camera reference frame $x_{c}$ using MeTRAbs-ACAE (37). We can rotate these keypoints into the global frame using $x_{n}=\hat{R}_{nc}x_{c}$ and define a loss function on the Euclidean distance in centimeters between keypoints from the model $\hat{x}_{n}$ with the video keypoints as:
@@
We defined implicit functions and computed the loss using Equinox (65) and JAX (66), optimizing with the Adam optimizer for 25,000 iterations of gradient descent. Weight decay of $1\mathrm{e}{-5}$ was applied to the implicit function parameters. With Optax, we set the learning rate to start at $1\mathrm{e}{-3}$ and decay exponentially to $1\mathrm{e}{-6}$ (67, 66).
For MMMC (26), we found random initialization of the network sufficient for optimizing implicit representations. However, for monocular videos, this often failed to converge if the subject wasn’t initially visible to the camera. To address this, we adjusted each trial’s implicit function by biasing the final layer to place the pelvis 1.5 meters from the camera and initializing $R_{nc}$ outputs to the median of observed $R_{nc}$ values.
@@
The GDI maps cycle-aligned joint angles to a lower-dimensional subspace in which we can measure distance from a normative population. We computed gait event timings for every self-selected gait trial using a pretrained transformer from our prior work (31), after discarding the first and last step. This generated 13,609 gait cycles from the PBL system. Of these, 1,267 were of control participants and we use these and 4,675 gait cycles from our data collection at the American Society of Biomechanics meeting and 3,066 cycles from our MMMC system as normative data.
@@
While the original GDI (42) used singular value decomposition to reduce this high dimensional data, more recent findings (43) demonstrate that principal component analysis (PCA) finds a more suitable subspace. We fit PCA to our 150 $\times$ 21,350 matrix, finding that 12 components accounted for 95% of the variance.
@@
To compare population-level GDIs, we first assessed normality using the Shapiro-Wilk test (68). For comparisons between two normally distributed populations, we used Student’s t-test. When at least one distribution was non-Gaussian, we used the Mann-Whitney U test to assess group differences.
@@
To assess the responsiveness of various measures to surgical intervention in the CM cohort, we computed the Standardized Response Mean (SRM) (45, 46), defined as:
@@
where ${D}_{x}$ is the vector of pre- to post-surgery differences in measure $x$, $\bar{D}_{x}$ is the mean of these differences, and $SD(D_{x})$ is their standard deviation. We calculated SRMs for the mJOA, cadence, double support time, and GDI to compare their sensitivity to change.
To estimate confidence intervals for the SRM, we employed the jackknife technique (47) as was done in (46). This nonparametric resampling method systematically recomputes the SRM by omitting one observation at a time, producing a set of jackknife replicates. These replicates are then used to compute pseudo-values, which approximate the influence of each observation on the overall SRM. The mean of these pseudo-values serves as a bias-corrected point estimate of the SRM, and their variability provides an estimate of its standard error. A t-distribution is then used to construct confidence intervals, enabling inference about the SRM’s precision without assuming normality of the underlying data.
@@
Overall, median joint angle error (MJAE) was approximately 3° across datasets, populations, and activities. Below, we highlight key trends and deviations from this general pattern, offering possible explanations and practical takeaways for users. However, we do not perform extensive statistical testing across the many variables examined, as the focus of this work is on enabling accessible motion capture rather than achieving the most precise measurement possible.
Table 1: Median Joint Angle Errors (median (nIQR)) for clinical and BML-MoVi datasets.
@@
Across datasets, the BML-MoVi dataset exhibited the lowest MJAE (2.74°; Table S1, Fig. 4C), which is expected since it included only control participants. The MMMC dataset showed similar errors when using a handheld camera (2.79°), but higher errors with a static camera (2.96°). This increase is likely due to the greater distance between the participant and camera in the static setup, whereas the handheld condition involved a researcher following the participant more closely (Fig. 2). Some of the added error in the MMMC dataset may also be attributable to noise in the MMMC system itself, despite the use of multiple cameras.
@@
When comparing clinical populations within the MMMC dataset, control participants showed the lowest MJAE (2.51°), while participants with neurological conditions showed the highest (3.32°; Table S2). This is consistent with expectations: many neurological participants used assistive devices—such as canes, walkers, or ankle-foot orthoses—which can obstruct joint visibility and hinder keypoint detection. Additionally, physical assistance from a therapist using a gait belt may contribute to further occlusion and error. Lower limb prosthesis users (LLPUs) also exhibited elevated MJAE (2.97°), reflecting known limitations of keypoint detectors when applied to prosthetic limbs (30).
@@
Across different activities, MJAE remained relatively stable (Fig. 4A). Unsurprisingly, standing had the lowest MJAE (1.80°), followed by the Four Square Step Test (FSST; 2.47°), both of which involve minimal occlusion. In contrast, activities such as the Timed Up and Go (3.17°), tandem gait (3.22°), and walking (2.95°) yielded slightly higher errors—likely due to partial leg occlusions caused by the oblique handheld filming angle used during much of these trials.
@@
Joint-level trends in MJAE were largely consistent across populations (Table S2). Notably, ankle MJAE exceeded 5° in all populations, indicating a clear need for improved foot keypoint detection. Elbow flexion error approached 5° in clinical cohorts, likely because one arm is frequently occluded during walking. Hip flexion error was higher than expected in the BML-MoVi and static-camera MMMC datasets (Table S1). These datasets featured more frontal views of the participant, which—as shown in our viewpoint analysis (Fig. 3C)—can increase sagittal plane error. Interestingly, this often manifested as a consistent offset rather than an error in the joint’s range of motion. A more detailed investigation of these effects, including marker placement and viewpoint-dependent bias in keypoint detection, is left for future work.
@@
Table 2: Median Joint Angle Error (median (nIQR)) for handheld smartphone reconstruction across different clinical populations.
Joint MJAE (nIQR)
@@
Across all datasets and modalities, 2D and 3D residuals between reconstructed and detected joint locations were generally low at $<$5 pixels for 2D reprojection errors and $<$1.5 cm for 3D marker errors. Correlations between fitting residuals and MJAE were significant but low: 0.35 and 0.24 for 2D and 3D residuals, respectively (Fig. S2A,C). While this suggests there may be some relationship between the closeness of our model fit to the actual joint angle accuracy, it is quite weak, at least at the levels most of our reconstructions achieved. This likely highlights the fact that the underlying computer vision algorithm does not have well calibrated confidence scores, something we aim to address in the future (49, 25). Importantly, we do not see any obvious issues with our approach generalizing to clinical settings as there were not large differences in either residual between in-lab and in-clinic cohorts (Fig. S2B,D).
Figure 2: Fitting Residuals. Correlation between 2D and 3D fitting residuals and MJAE is shown in (A), (C), respectively. 2D and 3D fitting residuals distributed across clinical populations is shown in (B) and (D), respectively.

### v2

@@ Acknowledgements.
The outputs of $f_{\phi}$ corresponding to rotations (i.e., not the pelvis location) are passed through a tanh nonlinearity to limit it to (1,1)(-1,1) followed by scaling this range to match the biomechanical model joint limits. Joint angles, $\theta(t)$, are then passed through the forward kinematic equation above (Eq. 1) to obtain the 3D joint locations, $\mathbf{x}(t)$, at that time.
@@
We define the global reference frame $\{n\}$ for our model output. Since the PBL system tracks changes in phone orientation from the smartphone IMUs, we are able to relate the orientation of the camera frame $\{c\}$ to the orientation of the world frame at any point in time using $R_{nc}\in SO(3)$. We follow the convention presented in [lynch_modern_2017] for rotations where $R_{ab}$ represents frame $\{b\}$ relative to frame $\{a\}$.
@@
Each trial is represented as a unique implicit function. We jointly optimize the parameters of the implicit representations for N trials $\{\phi_{0},\phi_{1},...,\phi_{N-1}\}$, the body scaling, and site offset parameters $\beta$, to follow 2D and 3D keypoints extracted from video as well as orientation data recorded from the smartphone. This approach of jointly learning kinematics and body scaling and marker offset for the entire session has been called bilevel optimization [werling_addbiomechanics_2023]. For example, a session consisting of 10 trials would jointly optimize 10 implicit functions $f_{\theta_{i}}$ and a single set of body scaling parameters $\beta$. A single trial required 1 minute to optimize on a NVIDIA A100 GPU and 4.5 minutes to optimize on a consumer-grade (NVIDIA GeForce RTX 3050 Laptop) GPU. Trial length has little effect on runtime, and additional trials add only an incremental processing cost.
[…]
@@
We obtain a pure video estimate of the 3D keypoint locations in the camera reference frame $x_{c}$ using MeTRAbs-ACAE [sarandi_learning_2023]. We can rotate these keypoints into the global frame using $x_{n}=\hat{R}_{nc}x_{c}$ and define a loss function on the Euclidean distance in centimeters between keypoints from the model $\hat{x}_{n}$ with the video keypoints as:
@@
We defined implicit functions and computed the loss using Equinox [kidger_equinox_2021] and JAX [deepmind2020jax], optimizing with the Adam optimizer for 25,000 iterations of gradient descent. Weight decay of $1\mathrm{e}{-5}$ was applied to the implicit function parameters. With Optax, we set the learning rate to start at $1\mathrm{e}{-3}$ and decay exponentially to $1\mathrm{e}{-6}$ [kingma_adam_2017, deepmind2020jax].
For MMMC [cotton_differentiable_2025], we found random initialization of the network sufficient for optimizing implicit representations. However, for monocular videos, this often failed to converge if the subject wasn’t initially visible to the camera. To address this, we adjusted each trial’s implicit function by biasing the final layer to place the pelvis 1.5 meters from the camera and initializing $R_{nc}$ outputs to the median of observed $R_{nc}$ values.
Validation Metrics
We evaluated normality of errors at two levels using the Shapiro–Wilk test [shapiro_analysis_1965]: (1) the distribution of joint angle errors within each trial and (2) the distribution of trial-level error summaries across trials. Both levels were non-Gaussian. Consequently, we report the Median Absolute Error, computed by first taking the median joint-angle error within each trial and then taking the median of these per-trial medians across trials. The Median Absolute Error includes the following degrees of freedom: hips (2×3), knees (2×1), ankles (2×1), lumbar spine (3), neck (3), shoulders (2×2), and elbows (2×1).
When aggregating results across trials, we also report the normalized interquartile range, $nIQR=IQR\cdot 0.7413$.
For comparison to prior work, including OpenCap, we additionally report the Mean Absolute Error following the procedure in [uhlrich_opencap_2023]. This Mean Absolute Error calculation includes the following joints: pelvis rotations (3), hips (2×3), knees (2×1), ankles (2×2), and lumbar (3).
To elucidate where in the gait cycle errors occur, we extracted 1,971 gait cycles from our MMMC dataset with paired MMMC (multi-camera) and PBL (single-camera) recordings and performed a paired SPM1D t-test [pataky_generalized_2010, pataky_one-dimensional_2012] between hip and knee flexion angles.
@@
The GDI maps cycle-aligned joint angles to a lower-dimensional subspace in which we can measure distance from a normative population. We computed gait event timings for every self-selected gait trial using a pretrained transformer from our prior work [cotton_transforming_2022], after discarding the first and last step. This generated 13,609 gait cycles from the PBL system. Of these, 1,267 were of control participants and we use these and 4,675 gait cycles from our data collection at the American Society of Biomechanics meeting and 3,066 cycles from our MMMC system as normative data.
@@
While the original GDI [schwartz_gait_2008] used singular value decomposition to reduce this high dimensional data, more recent findings [marks_measuring_2018] demonstrate that principal component analysis (PCA) finds a more suitable subspace. We fit PCA to our 150 $\times$ 21,350 matrix, finding that 12 components accounted for 95% of the variance.
@@
To compare population-level GDIs, we first assessed normality using the Shapiro-Wilk test [shapiro_analysis_1965]. For comparisons between two normally distributed populations, we used Student’s t-test. When at least one distribution was non-Gaussian, we used the Mann-Whitney U test to assess group differences.
@@
In addition to the Wilcoxon effect size calculated to assess responsiveness to surgical intervention, we computed the Standardized Response Mean (SRM) [husted_methods_2000, liang_comparisons_1990, werner_concurrent_2020], defined as:
@@
where ${D}_{x}$ is the vector of pre- to post-surgery differences in measure $x$, $\bar{D}_{x}$ is the mean of these differences, and $SD(D_{x})$ is their standard deviation. We calculated SRMs for the mJOA, cadence, double support time, and GDI.
To estimate confidence intervals for the SRM, we employed the jackknife technique [miller_jackknife-review_1974] as was done in [liang_comparisons_1990]. This nonparametric resampling method systematically recomputes the SRM by omitting one observation at a time, producing a set of jackknife replicates. These replicates are then used to compute pseudo-values, which approximate the influence of each observation on the overall SRM. The mean of these pseudo-values serves as a bias-corrected point estimate of the SRM, and their variability provides an estimate of its standard error. A t-distribution is then used to construct confidence intervals, enabling inference about the SRM’s precision without assuming normality of the underlying data.
@@
Overall, Median Absolute Error of joint angles was approximately 3° across datasets, populations, and activities. Below, we highlight key trends and deviations from this general pattern, offering possible explanations and practical takeaways for users. However, we do not perform extensive statistical testing across the many variables examined, as the focus of this work is on enabling accessible motion capture rather than achieving the most precise measurement possible.
Table 1: Median Absolute Error (median (nIQR)) for clinical and BML-MoVi datasets.
@@
Across datasets, the BML-MoVi dataset exhibited the lowest Median Absolute Error (2.74°; Table S1, Fig. 1C), which is expected since it included only control participants. The MMMC dataset showed similar errors when using a handheld camera (2.79°), but higher errors with a static camera (2.96°). This increase is likely due to the greater distance between the participant and camera in the static setup, whereas the handheld condition involved a researcher following the participant more closely (Fig. 3). Some of the added error in the MMMC dataset may also be attributable to noise in the MMMC system itself, despite the use of multiple cameras.
@@
When comparing clinical populations within the MMMC dataset, control participants showed the lowest Median Absolute Error (2.51°), while participants with neurological conditions showed the highest (3.32°; Table S2). This is consistent with expectations: many neurological participants used assistive devices—such as canes, walkers, or ankle-foot orthoses—which can obstruct joint visibility and hinder keypoint detection. Additionally, physical assistance from a therapist using a gait belt may contribute to further occlusion and error. Lower limb prosthesis users (LLPUs) also exhibited elevated Median Absolute Error (2.97°), reflecting known limitations of keypoint detectors when applied to prosthetic limbs [cimorelli_validation_2024].
Figure 1: Joint Angle Error Distributions. A) In the MMMC dataset, Median Absolute Error distributions remain relatively stable when aggregated across activity types, participant populations (B), as well as between handheld (moving) and static cameras. C) Joint-specific Median Absolute Error distributions from both the MMMC and BML-MoVi datasets.
@@
Across different activities, Median Absolute Error remained relatively stable (Fig. S1A). Unsurprisingly, standing had the lowest Median Absolute Error (1.80°), followed by the Four Square Step Test (FSST; 2.47°), both of which involve minimal occlusion. In contrast, activities such as the Timed Up and Go (3.17°), tandem gait (3.22°), and walking (2.95°) yielded slightly higher errors—likely due to partial leg occlusions caused by the oblique handheld filming angle used during much of these trials.
@@
Joint-level trends in Median Absolute Error were largely consistent across populations (Table S2). Notably, ankle Median Absolute Error exceeded 5° in all populations, indicating a clear need for improved foot keypoint detection. Elbow flexion error approached 5° in clinical cohorts, likely because one arm is frequently occluded during walking. Hip flexion error was higher than expected in the BML-MoVi and static-camera MMMC datasets (Table S1). These datasets featured more frontal views of the participant, which—as shown in our viewpoint analysis (Fig. 4C)—can increase sagittal plane error. Interestingly, this often manifested as a consistent offset rather than an error in the joint’s range of motion. A more detailed investigation of these effects, including marker placement and viewpoint-dependent bias in keypoint detection, is left for future work.
@@
Table 2: Median Absolute Error (median (nIQR)) for handheld smartphone reconstruction across different clinical populations.
Joint Median Absolute Error (nIQR)
@@
Table 3: Lower Body Mean Absolute Errors. Computed over trials (n = 1702) and participants (n = 107), and the reported mean is an average over activities and degrees of freedom (three for pelvis orientation, three for the lumbar, three per hip, one per knee, and two per ankle.)
(a) Across Activity
Activity FSST Standing TUG Tandem Walking All
Mean Absolute Error (deg) 4.14 (2.13) 2.19 (1.41) 5.17 (1.47) 4.91 (2.67) 4.46 (2.07) 4.18
(b) Across Population
Population Control LLPU Neurologic Pediatric All
Mean Absolute Error (deg) 3.90 (1.82) 4.29 (2.09) 4.49 (1.94) 4.11 (1.81) 4.20
Mean Absolute Error
Mean Absolute Errors calculated with the same joints as used in OpenCap [uhlrich_opencap_2023] were below 5 degrees for most activities and populations (Table S3). TUG had the highest Mean Absolute Error at 5.17°, and the neurologic population had the highest Mean Absolute Error at 4.49°. For the most faithful comparison to OpenCap, we computed Mean Absolute Error on the control population during walking, finding 4.32 (2.06)° of error, which approaches OpenCap’s reported error of 4.1° on the same population and activity.
SPM1D Comparison
Figure 2: SPM1D Comparison of Multi Camera and Monocular. Raw kinematic traces (n=1971) compared between the Multi-Camera (MMMMC) and Single Camera (PBL) modalities of the MMMC dataset are compared using paired SPM1D t-tests, revealing significant differences at almost every timepoint in the gait cycle. Dashed lines in second row represent p < 0.05 significance threshold.
Paired SPM1D t-tests on per-step kinematics showed significant differences in knee and hip angles across nearly the entire gait cycle (Fig. S2). Given the large sample size (n = 1,971), these results are expected, and the absolute differences remain small. Hip flexion was consistently underestimated by the single-camera method, suggesting a potential pelvis-tilt bias, possibly driven by viewpoint-dependent errors (Fig. 4C), as most recordings were captured from behind. Knee flexion was also underestimated during stance and slightly overestimated near peak flexion. Despite the statistical significance, the magnitude of these errors is modest.
Figure 3: Root Translation Error. Body translation in 3D space reconstructed from smartphone (green, solid) closely mirrored ground truth (grey, dashed). Note that our method accounts for smartphone rotation, but not translation.
Standardized Response Mean (SRM)
Table 4: Standardized Response Mean (SRM) due to Surgical Intervention SRM is a unitless measure of change in metric before and after surgery, with 95% confidence intervals.
Metric SRM
Patient-Reported mJOA 0.20 (-0.38 0.78)
mJOA-LE 0.14 (-0.45 0.72)
Smartphone-Based Cadence 0.34 (-0.20 0.88)
DST -0.48 (-0.88 -0.07)
GDI 0.45 (0.00 0.91)
SRM results (Table S4) aligned closely with the effect sizes reported in Table 2. For the smartphone-based metrics, SRMs for pre/post-operative change were 0.34 for cadence, –0.48 for DST, and 0.45 for GDI. Patient-reported outcomes showed smaller SRMs of 0.20 (mJOA) and 0.14 (mJOA-LE). Notably, the confidence intervals for DST and GDI did not cross zero, whereas all other intervals did.
@@
Across all datasets and modalities, 2D and 3D residuals between reconstructed and detected joint locations were generally low at <<5 pixels for 2D reprojection errors and <<1.5 cm for 3D marker errors. Correlations between fitting residuals and Median Absolute Error were significant but low: 0.35 and 0.24 for 2D and 3D residuals, respectively (Fig. S5A,C). While this suggests there may be some relationship between the closeness of our model fit to the actual joint angle accuracy, it is quite weak, at least at the levels most of our reconstructions achieved. This likely highlights the fact that the underlying computer vision algorithm does not have well calibrated confidence scores, something we aim to address in the future [pierzchlewicz_platypose_2024, cotton_biomechanical_2025]. Importantly, we do not see any obvious issues with our approach generalizing to clinical settings as there were not large differences in either residual between in-lab and in-clinic cohorts (Fig. S5B,D).
Figure 5: Fitting Residuals. Correlation between 2D and 3D fitting residuals and Median Absolute Error is shown in (A), (C), respectively. 2D and 3D fitting residuals distributed across clinical populations is shown in (B) and (D), respectively.